OpenStudy (anonymous):

Five times the difference of twice a number and nine is negative twenty-five

OpenStudy (anonymous):

Let x = the number 5(2x - 9) = -25 Simplifying this... 10x - 45 = -25 +45 +45 -------------- 10x = 20 /10 /10 -------------- x = 2 WHAT YOU DO TO ONE SIDE OF THE EQUATION, YOU MUST DO TO THE OTHER SIDE. Hope I helped ;)

OpenStudy (anonymous):

i see were i went wrong i put + 45 instead of -45

OpenStudy (anonymous):

yes u did thank u

OpenStudy (anonymous):

Your welcome, If I helped you, would you mind awarding me with a medal? thanks.